A force of 3000 N moves a mass of 500 Kg a distance of 8m in 15 seconds. How much work has been done?
1
Expert's answer
2019-03-21T10:45:20-0400

We can find the work done from the formula:


"W = Fs = 3000 N \cdot 8 m = 2.4 \cdot 10^4 J."

Answer:

"W = 2.4 \cdot 10^4 J."